# Second-order Optimization of Gaussian Splats with Importance Sampling

### v1

@@ 4.2 Matrix-free Preconditioned Conjugate Gradient (PCG) Solver for 3DGS
Although this optimizer is able to converge to the final solution in a very limited number of steps, it is 3×3$\times$3 × slower than our final method, as shown in Table 3. The main reason for this is that the conjugate gradient algorithm needs to run several iterations, therefore, we need to compute the $\mathbf{J}^{\top}\mathbf{J}\mathbf{v}$ product from Eq. 8 multiple times. This kind of computational complexity commonly arises when a second-order optimizer is used in large-scale numerical systems. The common approximations methods are diagonal [5, 28, 54, 47] and block-diagonal [32, 44, 6] approximation of the Hessian or $\mathbf{J^{\top}J}$ matrix.
@@
We observe that the Jacobians in 3DGS do not exhibit a dominant diagonal or block-diagonal structure, as illustrated in Fig. 3. This arises because each pixel is rendered through the interaction of multiple Gaussians, as described in Eq. 4. In other words, no single Gaussian independently represents a surface; instead, it relies on the collective contribution of surrounding Gaussians to accurately capture the ground truth. This leads to many off-diagonal and off-block-diagonal entries in $\mathbf{J^{\top}J}$. Therefore, the common diagonal and block-diagonal approximations do not provide good approximations. To this end, we propose to estimate the loss function with our proposed view and importance sampling.
@@
4.3 View Sampling
Second-order methods like LM are typically used in deterministic settings where the full objective is evaluated at every iteration. If this is not the case, estimating local curvature can be problematic and become unreliable [8]. This poses a challenge in the 3DGS setting, where the number of views can exceed the hundreds, as incorporating all of them at the same time is infeasible. Yet, to compute meaningful gradients, we must find an effective way to approximate the full normal equation in Eq. 8.
To this end, we introduce a view sampling approach that allows us to get a diverse set of views in each batch. We run K-Means clustering to partition the camera locations into batch size number of clusters (e.g., 8 clusters created for a batch size of 8). An image is then randomly selected from every cluster. This approach ensures that each batch captures a more balanced and diverse set of views, resulting in a more accurate estimation of the curvature information. As evidenced by Table 2, this method converges to higher scores compared to the random sampling of the cameras.
4.4 Estimation of Loss Function with Importance Sampling
Additionally, to improve the efficiency and feasibility of our second-order optimizer for 3DGS, we further exploit the sparsity property within the 3DGS representation, which has been discussed in Sec. 4.2. Yet, at the same time, as observed in Fig. 3, there does not exist a well-organized structure to the sparsity. Instead, in this paper we propose to adopt an importance sampling approach, where we focus more on the important elements that give meaningful gradients.
We begin our derivation by rewriting our least squares loss function in Eq. 5 as an integral. Then, the loss function is given as:
[…]
@@ 4.4 Estimation of Loss Function with Importance Sampling
[…]
@@
When we apply the Gauss-Newton approximation to Hessian matrix, the term with second derivative in Eq. 14 is ignored, and we are left with an estimation of $\mathbf{J^{\top}J}$.
@@
In this work, we experimented with simple uniform distribution $q_{\text{uniform}}(\mathbf{r})$ and a loss-based distribution $q_{\text{loss}}(\mathbf{r})$ to sample the pixels. Our intuition is that prioritizing pixels with high loss values facilitates the correction of the Gaussians responsible for these discrepancies. We define the likelihood of pixel $i$ using Softmax normalization as:
$q_{\text{loss}}(i)=Softmax(\|r_{i}\|)=\frac{\exp(\|r_{i}\|)}{\sum_{j}^{N}\exp(\|r_{j}\|)}$ (15)
Comparisons between the uniform and loss-based distribution are given in Table 3.
@@
In practice, sampling pixels over the image does not efficiently integrate with GPU programming and the backward pass of 3DGS. The reason is that an image is divided into $16\times 16$ tiles and when the pixels are selected randomly, some tiles get more samples than others, causing an unbalanced workload among thread blocks. Moreover, the random nature of the sampling does not allow fixed thread assignment per tile. Therefore, we use a stratified sampling strategy by distributing samples among tiles and perform importance sampling inside them. In other words, instead of sampling $N$ pixels from the entire image, we sample $N/\mathcal{T}$ pixels from each tile, where $\mathcal{T}$ is the total number of tiles in the image.
@@
The sampling mechanism allows us to achieve 3×3$\times$3 × speedup over the non-approximated LM optimizer, while maintaining similar performance, as shown in Table 3.
4.5 Learning Rate Scheduler
@@
When second-order optimizers are applied to large-scale systems, numerical errors can occur, causing the solution vector $\Delta\boldsymbol{\beta}$ to overshoot the true loss landscape. To determine the optimal learning rate, line search algorithms or trust region methods [2, 18, 37, 11, 41, 7] are usually employed. However, these methods require additional forward or backward passes, increasing the computational overhead and slowing down the overall algorithm. Rather than relying on these methods, we estimate the learning rate by constraining the maximum update of color values. In 3DGS, the color parameters range between $\approx-1.77$ and $\approx 1.77$ due to the level 00 spherical harmonic coefficient. This gives us a natural bound for the color parameters. We trust the update direction $\Delta\boldsymbol{\beta}$, as long as the resulting color change does not exceed 1. If the change surpasses this threshold, we scale $\Delta\boldsymbol{\beta}$ so that the maximum change remains 1. Notably, this scaling is applied uniformly across all parameters, regardless of their type (e.g., opacity, color, mean, scale, or rotation). The effectiveness of this learning rate heuristic is demonstrated in Table 4. Additionally, the supplementary document provides examples of learning rates assigned at each iteration.
Refer to caption Refer to caption Refer to caption
Figure 4: We plot the test set performance of baseline methods evaluated on the Lego scene. Our method achieves comparable metrics in significantly less time. All optimizers use the MSE loss function.
@@ 5 Results
Datasets and Metrics. We conduct our main experiments on the synthetic NeRF datasets [35], which include $8$ different scenes at $800\times 800$ resolution, each having 100 training and 200 testing camera views. To initialize the Gaussians, we follow 3DGS [21] and position $10{,}000$ Gaussians at random locations with random colors inside of the cube encapsulating the object. We also report results on Tanks &\&& Temples [24], which includes $2$ scenes: a truck, and a train scene. In these datasets, the Gaussians are initialized with structure-from-motion point clouds, and the number of Gaussians is $136{,}029$ and $182{,}686$, respectively. For all experiments, we use the default training and test splits. We report the test set performance based on structural similarity index (SSIM) [51], learned perceptual image patch similarity score (LPIPS) [56], and peak signal-to-noise ratio (PSNR).
Baselines. We compare our optimizer against Adam [23], RMSprop [50], and SGD with momentum [45], using their respective implementations in PyTorch. The momentum term of SGD and second-moment factor of RMSprop is set to $0.99$. We used the default learning rates given in 3DGS [21] for Adam and RMSProp optimizers. For SGD, we used the learning rates $0.16$ for the location, $0.2$ for color and $0.1$ for other parameters. We also apply a learning rate schedule for the mean value parameters, as suggested in 3DGS. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [30], which enhances the efficiency of the backward computation by storing the gradient state at every $32^{\text{nd}}$ splat. All baselines are run for $10{,}000$ iterations.
Implementation Details. All experiments are performed on a single NVIDIA Tesla A40 GPU. We use a uniform damping parameter $\lambda=0.01$ across all experiments. Our method, as well as all baselines, optimize using the mean squared error (MSE) loss function. We run our method for $200$ iterations, setting the batch size to $8$ and the number of sampled pixels per tile ($N$) to $32$, using our loss-based sampling distribution $q_{\text{loss}}$. Maximum conjugate gradient iterations start from $3$ and increase to $8$ after the $50^{th}$ iteration.
@@ 5.1 Comparison
Next, we evaluate our method, named as Levenberg-Marquardt with Importance Sampling (LM-IS), on synthetic NeRF datasets and present the results in Table 1. Our approach achieves strong performance while significantly reducing computation time. We report average metrics across all the synthetic NeRF datasets. Additionally, Figure 4 illustrates the test set performance on the Lego scene over time. As shown, our second-order optimizer converges rapidly and achieves competitive results across all metrics. Overall, our method demonstrates substantial improvements over the baselines in both efficiency and accuracy.
Another observation is that SGD with momentum consistently underperforms compared to other optimizers, despite our efforts at hyperparameter tuning. In fact, we had to truncate its training process to fit its results within the same plot as other optimizers. This highlights the importance of adaptive learning rates for optimizing 3DGS, a feature incorporated in all other evaluated optimizers. We also did not observe significant speed benefits from the optimizations of Taming 3DGS, which is likely due to a comparatively lower number of Gaussians in this setting that reduces the advantages of gradient caching.
Please refer to the supplemental document for the results using SSIM loss, as well as results on the Tanks &\&& Temples dataset.
Table 1: We report LPIPS, SSIM and PSNR scores averaged across NeRF synthetic test datasets. Our method achieves similar quality in a shorter amount of time.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
3DGS-SGD 0.331 0.488 18.56 164.2
3DGS-RMSprop 0.159 0.856 25.52 54.42
3DGS-Adam 0.162 0.855 25.49 58.46
Taming3DGS-Adam 0.147 0.862 26.01 60.73
LM-IS (Ours) 0.144 0.867 25.82 12.76
@@ 5.2 Ablation Studies
In this section, we ablate various design choices incorporated into our optimizer. All of the results in this section are obtained by averaging the metrics across all datasets.
@@
View Sampling. We propose a view sampling approach in Sec. 4.3 to ensure that our optimizer can effectively estimate the local curvature of the loss landscape while still maintaining a relatively low batch size, through sampling of diverse views in a batch. In Tab. 2, we ablate our design choice, where we observe that our method shows improvements as compared to random view sampling. Furthermore, we also report results on lower batch sizes, where the performance is significantly lower, further verifying that second-order optimizers indeed require larger batch sizes than first-order optimizers.
Table 2: Our optimizer relies on view sampling to obtain meaningful Jacobian approximations. If the number of images in the batch is limited, or if we use random sampling of images, the performance of the optimizer drops.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Cluster - Batch Size = 1 0.372 0.153 13.11
Cluster - Batch Size = 2 0.329 0.472 15.65
Cluster - Batch Size = 4 0.174 0.841 23.97
Random - Batch Size = 8 0.150 0.863 25.45
Cluster - Batch Size = 8 0.144 0.867 25.82
@@
Importance Sampling. Our proposed importance sampling approach in Sec. 4.4 is the core algorithm that provides significant speed benefits, enabling faster processing by effectively reducing the dimensionality of highly redundant Jacobians. Here, we compare our performance against three baselines in Tab. 3. Firstly, when no sampling is used at all, i.e., the full Jacobians are considered every time, the amount of training time required increases significantly, showing the importance of the sampling and approximation. When we adopt our importance sampling but a high number of sampled pixels per tile (e.g., $N=128$) is used, there are some gains over no sampling, but its efficiency is still suboptimal. Furthermore, when we adopt a simple random sampling ($q_{\text{uniform}}$) instead of our importance sampling, we also find that efficiency and accuracy drops. All these show the efficacy of our importance sampling method.
Table 3: This table illustrates the effectiveness of our loss-based sampling distribution. Even when the number of samples per tile ($N$) is low, we can effectively estimate the loss function and reach a similar performance that we would get without sampling. We also show that importance sampling performs slightly better than uniform sampling on average.
@@
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
LM - No Sampling 0.143 0.868 25.88 37.70
LM - $q_{\text{loss}}$ $N=128$ 0.142 0.868 25.91 24.20
LM - $q_{\text{uniform}}$ $N=32$ 0.147 0.866 25.68 13.16
LM - $q_{\text{loss}}$ $N=32$ 0.144 0.867 25.82 12.76
Learning Rate Heuristic. We compare our learning rate heuristic introduced in Sec. 4.5 against uniform learning rate, Armijo line search [2, 41], and a grid search method in Tab. 4.
@@
The grid search method selects the learning rate at each iteration by identifying the value that results in the greatest reduction in the objective function. Although this method is stable, the learning rates obtained are too pessimistic, and it cannot reach the quality of our heuristic. Armijo line search, or backtracking line search, improves the grid search method by picking the highest learning rate that results in sufficient reduction of the loss. While this approach has been successfully integrated as an improvement in deterministic second-order optimizers, our findings indicate that its effectiveness is limited in stochastic settings, such as ours, where optimization relies on multiple approximations. The performance of the line search methods can be tuned by using fewer candidate learning rates, however, we found that even with a large set of candidates, they fail to reach the quality of our heuristic within the same number of iterations.
Table 4: We compare the effect of various learning rate schedules. Our learning rate heuristic is able to assign dynamic learning rates resulting in rapid convergence rate and stability.
@@
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
Uniform $LR=0.5$ 0.504 0.365 9.46 18.32
Uniform $LR=0.3$ 0.380 0.544 13.64 15.57
Uniform $LR=0.1$ 0.161 0.855 25.37 14.33
Grid Line Search 0.262 0.756 21.64 761.7
Armijo Line Search 0.363 0.658 13.34 51.75
Our Heuristic (Sec. 4.5) 0.144 0.867 25.82 12.76

### v2

@@ 4.2 Matrix-free Preconditioned Conjugate Gradient (PCG) Solver for 3DGS
Although this optimizer is able to converge to the final solution in a limited number of steps, it is $4\times$ slower than our final method, as shown in Tab. 3. The main reason for this is that the conjugate gradient algorithm needs to run several iterations, therefore, we need to repeatedly compute the $\mathbf{J}^{\top}\mathbf{J}\mathbf{p}$ product arising on line $10$ of Alg. 1. This kind of computational complexity commonly arises when a second-order optimizer is used in large-scale numerical systems, and usually, an approximation method is used for efficiency. The common approximation methods are diagonal [becker1989improving, sophia, adahessian, pesky] and block-diagonal [kfac, roux2007topmoumoute, botev2017practical] approximation of the Hessian or $\mathbf{J^{\top}J}$ matrix.
@@
We observe that the Hessian approximation $\mathbf{J^{\top}J}$ in 3DGS does not exhibit a dominant diagonal or block-diagonal structure, as illustrated in Fig. 3. This arises because each pixel is rendered through the interaction of multiple Gaussians, as described in Eq. 4. In other words, no single Gaussian independently represents a surface; instead, it relies on the collective contribution of surrounding Gaussians to accurately capture the ground truth. This leads to many off-diagonal and off-block-diagonal entries in $\mathbf{J^{\top}J}$. Therefore, the common diagonal and block-diagonal approximations do not provide good approximations. To this end, we propose to estimate the loss function with our proposed view and residual sampling. The sampling techniques enable the matrix-free implementation as the cost of solving Eq. 8 is drastically reduced.
4.3 View Sampling
Second-order methods like LM are typically used in deterministic settings where the full objective is evaluated at every iteration. If this is not the case, estimating local curvature can be problematic and become unreliable [byrd2011use]. This poses a challenge in the 3DGS setting, where the number of views can exceed hundreds, as incorporating all of them at the same time is infeasible. Yet, to compute meaningful gradients, we must find an effective way to approximate the full normal equation in Eq. 8.
To this end, we introduce a camera sampling approach that allows us to get a diverse set of views in each batch. Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ be the set of all cameras. For each camera $c_{i}$, we form the feature vector:
$f_{i}=[\tilde{x}_{i},\tilde{y}_{i},\tilde{z}_{i},\;{dx}_{i},{dy}_{i},{dz}_{i}]$ (10)
[…]
4.4 Estimation of Loss Function with Residual Sampling
@@
Additionally, to improve the efficiency and feasibility of our second-order optimizer for 3DGS, we further exploit the sparsity property within the 3DGS representation, which has been discussed in Sec. 4.2. Yet, at the same time, as observed in Fig. 3, there does not exist a well-organized structure to the sparsity. Instead, in this paper, we propose to adopt a residual sampling approach.
We begin our derivation by rewriting our least squares loss function in Eq. 5 as an integral: Then, the loss function is given as:
$L=\int_{\Omega}r(\mathbf{p})^{2}\,d\mathbf{p}=\int_{\Omega}\ell(\mathbf{p})\,d\mathbf{p},$ (11)
@@
$L=\int_{\Omega}\ell(\mathbf{p})\frac{q(\mathbf{p})}{q(\mathbf{p})}\,d\mathbf{p}=\mathbb{E}_{\mathbf{p}\sim q(\mathbf{p})}\left[\frac{\ell(\mathbf{p})}{q(\mathbf{p})}\right].$ (12)
@@
When we apply the Gauss-Newton approximation to the Hessian matrix, the term with second derivative in Eq. 15 is ignored, and we are left with an estimation of $\mathbf{J^{\top}J}$.
@@
In this work, we evaluated a range of candidate distributions $q$ and found that simple uniform distribution works well. For a comparison of alternative distributions and their respective results, please refer to the Supplementary.
@@
In practice, sampling pixels over the image does not efficiently integrate with GPU programming and the backward pass of 3DGS. The reason is that an image is divided into $16\times 16$ tiles and when the pixels are selected randomly, some tiles get more samples than others, causing an unbalanced workload among thread blocks. Moreover, the random nature of the sampling does not allow fixed thread assignment per tile. Therefore, we use a stratified sampling strategy by distributing samples among tiles, and we perform sampling inside them. In other words, instead of sampling $N$ pixels from the entire image, we sample $N/\mathcal{T}$ pixels from each tile, where $\mathcal{T}$ is the total number of tiles in the image.
@@
The sampling mechanism allows us to achieve $4\times$ speedup over the non-approximated LM optimizer, while maintaining similar performance, as shown in Tab. 3.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s) Memory (GB)
3DGS-SGD 0.331 0.488 18.56 164 2.43
3DGS-RMSprop 0.159 0.856 25.52 54 2.43
3DGS-Adam 0.162 0.855 25.49 58 2.43
Taming3DGS-Adam 0.147 0.862 26.01 60 2.54
LM-RS (Ours) 0.139 0.872 26.14 13 3.07
Figure 4: The averaged results of competing methods on the synthetic NeRF dataset are presented. Our method achieves similar quality in a shorter amount of time. We also provide convergence plots on the Lego scene.
4.5 Dynamic Learning Rate Scheduler
@@
When second-order optimizers are applied to large-scale systems, numerical errors can occur, causing the solution vector $\Delta\boldsymbol{\beta}$ to overshoot the true loss landscape. To determine the optimal learning rate, line search algorithms or trust region methods [armijo1966minimization, hollein20243dgs, more1994line, conn2000trust, nocedal1999numerical, bottou2018optimization] are usually employed. However, these methods require additional forward or backward passes, increasing the computational overhead and slowing down the overall algorithm. Rather than relying on these methods, we estimate the learning rate by constraining the maximum update of color values. In 3DGS, the color parameters range between $\approx-1.77$ and $\approx 1.77$ due to the level 0 spherical harmonic coefficient. This gives us a natural bound for the color parameters. We trust the update direction $\Delta\boldsymbol{\beta}$, as long as the resulting color change does not exceed 1. If the change surpasses this threshold, we scale $\Delta\boldsymbol{\beta}$ so that the maximum change remains 1. Notably, this scaling is applied uniformly across all parameters, regardless of their type (e.g., opacity, color, mean, scale, or rotation). The effectiveness of this learning rate heuristic is demonstrated in Tab. 4, and examples of assigned learning rates are given in the Supplementary.
Figure 5: We evaluate the performance of the optimizers in the real-world scenes. Our optimizer attains better quality in shorter training time.
@@ 5 Results
Datasets and Metrics. We conduct our main experiments on the synthetic NeRF [nerf] and real-world Mip-NeRF360 indoor datasets [barron2021mipnerf]. In the synthetic dataset, we position 10,00010{,}000 Gaussians at random locations with random colors inside of the cube encapsulating the object. In the real-world scenes, the Gaussians are initialized with structure-from-motion point clouds as in 3DGS [3dgs]. For all experiments, we use the default training and test splits. We report the test set performance based on structural similarity index (SSIM) [ssim], learned perceptual image patch similarity score (LPIPS) [lpips], and peak signal-to-noise ratio (PSNR).
Baselines. In the synthetic dataset, we compare our optimizer against Adam [adam], RMSprop [rmsprop], and SGD with momentum [sgd_momentum], using their respective implementations in PyTorch. The gradients are computed using either the vanilla 3DGS or Taming 3DGS [taming], which enhances the efficiency of the backward computation by storing the gradient state at every $32^{\text{nd}}$ splat.
In the real-world scenes, we compare against vanilla 3DGS [3dgs] and 3DGS-LM [hollein20243dgs]. Since 3DGS-LM’s code is not publicly available, we contacted the authors, who kindly provided PSNR scores, runtime, and average memory usage across several datasets. In addition, they provided results obtained without the Adam optimizer, which are denoted in the tables as 3DGS-LM (w/o Adam). First-order baselines are run for $10,000$ iterations, and we performed a hyperparameter search to find the best learning rates.
Implementation Details. We conducted synthetic dataset experiments on a NVIDIA Tesla A40 GPU and real-world experiments on a single NVIDIA Tesla A100 GPU. Our method and the baselines optimize the mean squared error (MSE) loss. We set the sampled pixels per tile ($N$) to 32 in each experiment. We use a fixed damping parameter $\lambda$ of $0.1$ for the synthetic NeRF scenes and Mip-NeRF360 dataset. For the synthetic dataset, the batch size is fixed at $8$; for real-world datasets, it is initialized at $16$ and increased to $32$ after the 50th iteration. The maximum number of conjugate gradient iterations starts at $3$ for synthetic datasets and at $5$ for real-world datasets, rising to $8$ after the 50th iteration.
Table 1: This table presents the comparisons of competing methods averaged in the Mip-NeRF360 indoor dataset.
[…]
3DGS 0.854 28.21 252 5.81
3DGS-LM (w/o Adam) - 26.72 1186 $\approx$ 50
3DGS-LM - 28.31 398 $\approx$ 50
LM-RS (Ours) 0.859 28.59 190 14.75
@@ 5.1 Comparison
Next, we evaluate our method, named Levenberg-Marquardt with Residual Sampling (LM-RS), on synthetic NeRF scenes and present the results in Fig. 4. As shown, our second-order optimizer converges rapidly and achieves a substantial $\approx 5\times$ speedup over Adam. We observed that SGD with momentum consistently underperforms compared to other optimizers, despite our efforts at hyperparameter tuning. This highlights the importance of adaptive learning rates for optimizing 3DGS, a feature incorporated in all other evaluated optimizers. We also did not observe significant speed benefits from the optimizations of Taming 3DGS, which is likely due to a comparatively lower number of Gaussians in this setting that reduces the advantages of gradient caching.
In addition, we provide results in Mip-NeRF360 dataset in Tab. 1 and Fig. 5. Our method achieves $1.3\times$ speedup over Adam, $2\times$ over 3DGS-LM and $6\times$ speedup over 3DGS-LM (w/o Adam). Thanks to our matrix-free algorithm, we do not rely on the Adam optimizer at all, and our memory requirement is $3.5\times$ less than 3DGS-LM.
Please refer to the Supplementary for the results using SSIM loss, as well as quantitative results in Tanks & Temples [knapitsch2017tanks] and DeepBlending scenes [hedman2018deep].
@@ 5.2 Ablation Studies
In this section, we ablate various design choices incorporated into our optimizer.
@@
View Sampling. We propose a view sampling approach in Sec. 4.3 to ensure that our optimizer can effectively estimate the local curvature of the loss landscape while still maintaining a relatively low batch size, through sampling of diverse views in a batch. In Tab. 2, we ablate our design choice, where we observe that our method shows improvements compared to random view sampling. We also report performance with reduced batch size, where we observe that the performance is significantly lower.
Table 2: The table demonstrates the significance of high batch sizes and the effectiveness of our clustering-based camera sampler compared to random sampler. The results are averaged over Tanks & Temples scenes. (MBS= Maximum Batch Size)
View Sampler LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Clustering-Based, MBS=8 0.363 0.640 19.18
Clustering-Based, MBS=16 0.271 0.719 21.37
Random, MBS=32 0.276 0.722 21.68
Clustering-Based, MBS=32 0.253 0.731 21.83
@@
Residual Sampling Size. Our proposed sampling approach in Sec. 4.4 is the core algorithm that provides significant speed benefits, enabling faster processing by effectively reducing the dimensionality of highly redundant Jacobians. Here, we compare our performance against three baselines in Tab. 3. Firstly, when no sampling is used at all, i.e., the full Jacobians are considered every time, the amount of training time required increases significantly, showing the importance of sampling. When we adopt our pixel sampling, but use a high number of sampled pixels per tile (e.g., $N=128$, $N=64$), there are some gains over no sampling, but its efficiency is still suboptimal. All these show the efficacy of our residual sampling method.
Table 3: We show that sampling fewer pixels per tile (N) has little impact on quantitative performance while substantially increasing efficiency. The results are averaged over the MipNeRF-360 indoor scenes.
@@
Residual Sampler LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
LM - No Sampling 0.143 0.868 28.68 810
LM-RS - $N=128$ 0.159 0.860 28.63 470
LM-RS - $N=64$ 0.160 0.860 28.66 286
LM-RS - $N=32$ 0.161 0.859 28.59 190
@@
Dynamic Learning Rate Scheduler. We compare our learning rate heuristic introduced in Sec. 4.5 against uniform learning rate, Armijo line search [armijo1966minimization, nocedal1999numerical], and a grid search method in Tab. 4. The grid search method selects the learning rate at each iteration by identifying the value that results in the greatest reduction in the objective function. Although this method is stable, the learning rates obtained are too pessimistic, and it cannot reach the quality of our heuristic. Armijo line search improves the grid search method by picking the highest learning rate that results in sufficient reduction of the loss. While this approach has been successfully integrated as an improvement in deterministic second-order optimizers, our findings indicate that its effectiveness is limited in stochastic settings, such as ours, where optimization relies on multiple approximations. We also found that the uniform learning rate can diverge (LR=0.1) or behave suboptimally (LR=0.07) compared to our scheduler.
Table 4: We show that our dynamic learning rate algorithm can converge faster, compared to uniform learning rate or search methods. The results are averaged across Mip-NeRF360 indoor scenes.
@@
Learning Rate LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Time (s)
Uniform $LR=0.07$ 0.202 0.842 27.91 188
Uniform $LR=0.1$ 0.281 0.669 23.20 188
Grid Line Search 0.224 0.821 27.12 397
Armijo Line Search 0.264 0.763 23.66 416
Our Scheduler (Sec. 4.5) 0.161 0.859 28.59 190